Two charged particles are a distance of 1.82 m from each other. One of the particles has a charge of 6.89 nC, and the other has a charge of 4.18 nC.
(a)
What is the magnitude (in N) of the electric force that one particle exerts on the other?
1
Expert's answer
2020-10-07T10:02:19-0400

Here,

Q1 = 6.89nC

Q2 = 4.18nC

r12 = 1.82m

Force exerted by Q2 on Q1 is given by-

F12 = "\frac{kQ_1Q_2}{r_{12}^2}"

Here, k = "\frac{1}{4\pi \epsilon_0}" = 9"\times" 109

So,

F12 = "\frac{(9\times10^9)(6.89\times10^{-9})(4.18\times10^{-9})}{(1.82)^2}"

= 7.825 "\times10^{-8}" N

Since both charges are negative, the force between two charges will be repulsive in nature.